how can we find the minimum length of a line ,which lies between the to coordinate axes and passes through the point (a,b)? and what is the minimum length?

Grade:12th pass

1 Answers

Abhyudaya Mishra
12 years ago

use (x-x1)/cosØ=(y-y1)/sinØ=r